# BanditSpec: Adaptive Speculative Decoding via Bandit Algorithms

### v1

@@ Abstract
Speculative decoding has emerged as a popular method to accelerate the inference of Large Language Models (LLMs) while retaining their superior text generation performance. Previous methods either adopt a fixed speculative decoding configuration regardless of the prefix tokens, or train draft models in an offline or online manner to align them with the context. This paper proposes a training-free online learning framework to adaptively choose the configuration of the hyperparameters for speculative decoding as text is being generated. We first formulate this hyperparameter selection problem as a Multi-Armed Bandit problem and provide a general speculative decoding framework BanditSpec. Furthermore, two bandit-based hyperparameter selection algorithms, UCBSpec and EXP3Spec, are designed and analyzed in terms of a novel quantity, the stopping time regret. We upper bound this regret under both stochastic and adversarial reward settings. By deriving an information-theoretic impossibility result, it is shown that the regret performance of UCBSpec is optimal up to universal constants. Finally, extensive empirical experiments with LLaMA3 and Qwen2 demonstrate that our algorithms are effective compared to existing methods, and the throughput is close to the oracle best hyperparameter in simulated real-life LLM serving scenarios with diverse input prompts.
@@ 1 Introduction
A Large Language Model (LLM) is trained to predict the probability of the next token conditioned on all previous tokens (Brown et al., 2020; Touvron et al., 2023). This autoregressive decoding approach involves multiple forward passes, with each pass generating one token sequentially. Consequently, this process can lead to significant latency during inference.
Speculative decoding was introduced by Leviathan et al. (2023); Chen et al. (2023) to accelerate the inference of LLMs. The standard speculative decoding framework has been extended with improved performance since then. A thorough overview is presented at Appendix A. While the existing speculative decoding methods are diverse, most previous works adopt a fixed one across tasks, severely limiting their potential. For instance, when dealing with code debugging or grammar-checking tasks, the generated tokens are expected to resemble most of the input tokens. Therefore, retrieval-based speculative decoding techniques are preferred (Hu et al., 2024). In contrast, for story generation tasks, we expect the generated tokens to be more creative. Thus, a draft model with a high-temperature parameter is preferred over retrieval-based methods. The potential of these speculative decoding methods can only be exploited when the configuration of hyperparameters is well-aligned to the given task. There are existing works that attempt to achieve this goal, e.g., Zhou et al. (2024) distills the draft model during inference. Furthermore, even when the choice of the draft model is optimized, the associated hyperparameters can still be refined. For instance, Liu et al. (2024) and Huang et al. (2024) aim to optimize the speculation length in a training and training-free manner. Based on these observations, we ask the questions (see Figure 1) : given prefix prompts and candidate configurations of hyperparameters, is there a theoretically sound framework to model and solve the hyperparameters selection problem? Is there any training-free method that can adaptively choose the hyperparameters such that the latency of speculative decoding can be minimized?
@@ 2 Preliminaries
LLM Decoding We denote an LLM as $P:\mathcal{X}^{*}\rightarrow\bm{\Delta}_{\mathcal{X}}$, where $\mathcal{X}$ and $\mathcal{X}^{*}$ are the space of tokens and the space of all token sequences, respectively. Most LLMs predict this conditional probability via predicting the logits of the next token. Concretely, the LLMs predict $\log P(x_{t}\,|\,x_{1:t-1})$, where $x_{t}\in\mathcal{X}$ and $x_{1:t-1}\in\mathcal{X}^{*}$ are respectively the $t$-th token and first $t-1$ tokens. In the inference stage, LLMs use an additional temperature parameter $\gamma>0$ to predict the next token’s probability as $\texttt{softmax}(\gamma^{-1}\log P(\cdot\,|\,x_{1:t-1}))$, where softmax is the softmax operator. The results in our work hold for any $\gamma>0$, and we just denote $\texttt{softmax}(\gamma^{-1}\log P)$ as $P$ for ease of notation. When $\gamma>0$, we sample the next token from the predicted distribution, which is called sampling (sampling decoding). When $\gamma\downarrow 0$, the next token will be the token that corresponds to the highest logit value; this is called greedy decoding. We note that the greedy decoding is deterministic, i.e., the token is sampled from a degenerate distribution. These two families of decoding methods can be unified as Algorithm 1, where LLMs autoregressively generate tokens until the $\mathrm{EOS}$ token.
@@
Speculative Decoding As shown in Algorithm 1, the autoregressive decoding feature requires multiple forward inferences of LLMs $P$ sequentially. To reduce the number of forward inferences, Leviathan et al. (2023); Chen et al. (2023) proposed the vanilla speculative decoding algorithm which implements a draft model $Q$ to generate draft tokens and let the target model $P$ verify them in parallel. For completeness, we present and describe the vanilla speculative decoding algorithm in Appendix C.1. This vanilla specualtive decoding is then extended by some existing works, e.g., Miao et al. (2023); Cai et al. (2024) organizes the draft tokens as a tree, which improves the number of accepted tokens. The speculative decoding algorithm contains many hyperparameters, e.g., the draft model $Q$, and the tree structure in Miao et al. (2023); Cai et al. (2024). Most existing works keep these hyperparameters fixed for all the tasks. Some other works optimize the draft model in an online or offline manner (Liu et al., 2023) and the size of the tree (Chen et al., 2024), which are designed for specific considerations. In contrast, our work aims to derive a unified online hyperparameter selection algorithm that can be applied for any type of hyperparameters.
Multi-Armed Bandits The Multi-Armed Bandit (MAB) is a fundamental online decision-making problem (see Algorithm 7 for its dynamics). In its classical stochastic form, an agent chooses from $K$ arms, each of which delivers a reward sampled i.i.d. from an unknown but fixed distribution when pulled (Lattimore & Szepesvári, 2020). The goal is to select arms over $T$ rounds to maximize cumulative rewards. Two primary classes of algorithms—UCB-type (Auer et al., 2002b) and sampling-based methods (Russo et al., 2017)—have been developed and proven optimal in this setting. In the adversarial formulation, there are no assumptions on the reward distributions; rewards can evolve arbitrarily over time and may be correlated across arms (Auer et al., 2002a). Several algorithms, such as EXP3 and EXP4 (Auer et al., 2002a), are known to achieve optimal performance under these conditions. In this work, we frame the hyperparameter selection problem as an MAB problem and develop algorithms tailored to both stochastic and adversarial settings.
Notations: Let $[N]:=\{1,\cdots,N\}$. For a finite set $\mathcal{X}$, we denote the set of distributions supported on it as $\bm{\Delta}_{\mathcal{X}}=\{P:\mathcal{X}\rightarrow[0,1]\,|\,\sum_{x\in\mathcal{X}}P(x)=1,P(x)\geq 0\text{ for all }x\in\mathcal{X}\}$. The space of all finite length sequences whose components belong to $\mathcal{X}$ is denoted as $\mathcal{X}^{*}$, and we use $x_{1:L}\in\mathcal{X}^{L}\subseteq\mathcal{X}^{*}$ to denote a length-$L$ sequence. The Kullback–Leibler ($\mathrm{KL}$) divergence between two distributions $P$ and $Q$ is denoted as $\mathrm{KL}(P,Q)$.
@@ 3 Bandits for Adaptive Speculative Decoding
Algorithm 2 Specuative Decoding Subroutine (SpecDecSub)
@@
To put the problem in a mathematically sound way, we first specify a general speculative decoding subroutine (SpecDecSub) in Algorithm 2. The input of this subroutine is a prompt $\mathrm{pt}\in\mathcal{X}^{*}$, a target model $P$, a specification of hyperparameters $S$, and the maximum speculation length $L$, and the output is the accepted token sequence $x_{1:\tau}\in\mathcal{X}^{*}$. We provide two examples of the hyperparameter sets here. (1) If we adopt the vanilla speculative decoding (Algorithm 6) as Line 1, $S$ can be different draft models $Q:\mathcal{X}^{*}\rightarrow\bm{\Delta}_{\mathcal{X}}$, and $\mathcal{S}$ is the set of all the provided draft models. We would like to choose a draft model according to its training context, e.g. math, creative writing, to decode the current prefix. Then the problem we consider is how to adaptively select a proper draft model for speculative decoding via bandit algorithms. (2) If we adopt Medusa (Cai et al., 2024) as Line 1, $S$ can be different tree structures, and $\mathcal{S}$ is the set of plausible tree structures. In this problem, we would like to adaptively adjust the speculation tree structure according to the context.
@@
Before going to the algorithm design and the theoretical analysis, we would like to specify some important properties that are shared for any arm selection algorithm $\mathtt{ALG}$ and clarify the intuitions about our theoretical analysis. We denote the stopping time of the canonical decoding (Algorithm 1) and the generated sequence as $\tau_{\mathrm{c}}$ and $\mathrm{pt}_{\tau_{\mathrm{c}}}$, respectively.
@@
[…]
@@
Interpretation of the Desired Result. Given a prompt $\mathrm{pt}$ and bandit configuration $\nu$, BanditSpec adpatively selects the hyperparameter via $\mathtt{ALG}$ and learns the context. The stopping time regret (1) measures how the stopping time of $\textsc{BanditSpec}(\mathtt{ALG})$ compared to that of the (agnostic) best one $\textsc{BanditSpec}(\mathtt{ALG}_{i^{*}})$. By minimizing $\mathrm{Reg}(\mathtt{ALG},\mathrm{pt},\nu)$, we want to devise an $\mathtt{ALG}$ to (approximately) match the performance of $\mathtt{ALG}_{i^{*}}$. In particular, if $\mathrm{Reg}(\mathtt{ALG},\mathrm{pt},\nu)=o(\mathbb{E}[\mathrm{len}(\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})})])$, it implies that $\textsc{BanditSpec}(\mathtt{ALG})$ requires the same number of speculative decoding rounds as $\textsc{BanditSpec}(\mathtt{ALG}_{i^{*}})$ asymptotically even though the information about $S_{i^{*}}$ is not revealed at the beginning. In order words, $\textsc{BanditSpec}(\mathtt{ALG})$ learns the identity of $S_{i^{*}}$ quickly and the price for this learning process can be amortized over time. Additionally, when $\textsc{BanditSpec}(\mathtt{ALG})$ is deployed over diverse prompt inputs, we expect a significant acceleration of token generation compared to any fixed single speculative decoding method.
@@ 4 Modeling Tokens Stochastically
Figure 2: Illustration of our bandit model for choosing configurations to decode the next token, where UCB and EXP3 refer to UCBSpec and EXP3Spec, repectively.
@@ 4.1 Upper Bounds for the Stochastic Case
[…]
This aggregated algorithm, UCBSpec, is adapted from the classical UCB-1 algorithm in Auer et al. (2002b). The main differences are the confidence radius design $\mathrm{cr}_{i,t}$ and the stopping rule. We highlight that the form of $\mathrm{cr}_{i,t}$ is designed to fit the weak assumption of the number of accepted tokens. In fact, the proof of the regret of UCB-1 assumes that the values of each arm are generated before the pull of arms (Auer et al., 2002b; Lattimore & Szepesvári, 2020), which bifurcates from practical LLM inference scenarios. In contrast, we remove this strong restriction. The stopping rule of UCBSpec makes the analysis of our algorithm rather different from that of UCB-1. The stochasticity of the total number of arm pulls requires a novel regret decomposition analysis that is not presented in previous works.
@@ 4.2 Lower Bound for the Stochastic Case
This TGD was considered in the seminal works on speculative decoding (Leviathan et al., 2023; Chen et al., 2023).
@@
For the tightness of our algorithm, according to Note 15.3 in Lattimore & Szepesvári (2020), $\mathrm{kl}_{i}=O(\Delta_{i}^{2})$ when $\Delta_{i}$ is small. This indicates the dominating terms in the upper bound in Theorem 4.3 match the lower bound in Theorem 4.4 up to (possibly instance-dependent) constants. Futherfmore, because the truncated geometric distribution is more close to a sub-exponential family distribution, especially when $L$ is large, bandit algorithms built upon UCB1 (Auer et al., 2002b) are generally loose in some factors. In order to close the gap between the upper and lower bounds completely, KL-UCB (Garivier & Cappé, 2011) can possibly be adapted to this problem out of theoretical interest. However, on the practical side, KL-UCB demands solving an optimization problem at each round, which can be time-consuming during implementations. Thus, it does not perfectly align with our ultimate goal of LLM inference acceleration.

### v2

@@ Abstract
Speculative decoding has emerged as a popular method to accelerate the inference of Large Language Models (LLMs) while retaining their superior text generation performance. Previous methods either adopt a fixed speculative decoding configuration regardless of the prefix tokens or train draft models in an offline or online manner to align them with the context. This paper proposes a training-free online learning framework to adaptively choose the configuration of the hyperparameters for speculative decoding as text is being generated. We first formulate this hyperparameter selection problem as a Multi-Armed Bandit problem and provide a general speculative decoding framework BanditSpec. Furthermore, two bandit-based hyperparameter selection algorithms, UCBSpec and EXP3Spec, are designed and analyzed in terms of a novel quantity, the stopping time regret. We upper bound this regret under both stochastic and adversarial reward settings. By deriving an information-theoretic impossibility result, it is shown that the regret performance of UCBSpec is optimal up to universal constants. Finally, extensive empirical experiments with LLaMA3 and Qwen2 demonstrate that our algorithms are effective compared to existing methods, and the throughput is close to the oracle best hyperparameter in simulated real-life LLM serving scenarios with diverse input prompts.
@@ 1 Introduction
A Large Language Model (LLM) is trained to predict the probability of the next token conditioned on all previous tokens (Brown2020LanguageMA; Touvron2023Llama2O). This autoregressive decoding approach involves multiple forward passes, with each pass generating one token sequentially. Consequently, this process can lead to significant latency during inference.
Speculative decoding was introduced by leviathan2023fast; chen2023accelerating to accelerate the inference of LLMs. The standard speculative decoding framework has been extended with improved performance since then. A thorough overview is presented at Appendix A. While the existing speculative decoding methods are diverse, most previous works adopt a fixed one across tasks, severely limiting their potential. For instance, when dealing with code debugging or grammar-checking tasks, the generated tokens are expected to resemble most of the input tokens. Therefore, retrieval-based speculative decoding techniques are preferred (hu2024sam). In contrast, for story generation tasks, we expect the generated tokens to be more creative. Thus, a draft model with a high-temperature parameter is preferred over retrieval-based methods. The potential of these speculative decoding methods can only be exploited when the configuration of hyperparameters is well-aligned to the given task. There are existing works that attempt to achieve this goal, e.g., zhou2024distillspec distills the draft model during inference. Furthermore, even when the choice of the draft model is optimized, the associated hyperparameters can still be refined. For instance, liu2024optimizing and huang2024specdec++ aim to optimize the speculation length in a training and training-free manner. Based on these observations, we ask the questions (see Figure 1) : given prefix prompts and candidate configurations of hyperparameters, is there a theoretically sound framework to model and solve the hyperparameters selection problem? Is there any training-free method that can adaptively choose the hyperparameters such that the latency of speculative decoding can be minimized?
@@ 2 Preliminaries
LLM Decoding We denote an LLM as $P:\mathcal{X}^{*}\rightarrow\bm{\Delta}_{\mathcal{X}}$, where $\mathcal{X}$ and $\mathcal{X}^{*}$ are the space of tokens and the space of all token sequences, respectively. Most LLMs predict this conditional probability via predicting the logits of the next token. Concretely, the LLMs predict $\log P(x_{t}\,|\,x_{1:t-1})$, where $x_{t}\in\mathcal{X}$ and $x_{1:t-1}\in\mathcal{X}^{*}$ are respectively the $t$-th token and first t1t-1 tokens. In the inference stage, LLMs use an additional temperature parameter $\gamma>0$ to predict the next token’s probability as $\texttt{softmax}(\gamma^{-1}\log P(\cdot\,|\,x_{1:t-1}))$, where softmax is the softmax operator. The results in our work hold for any $\gamma>0$, and we just denote $\texttt{softmax}(\gamma^{-1}\log P)$ as $P$ for ease of notation. When $\gamma>0$, we sample the next token from the predicted distribution, which is called sampling (sampling decoding). When $\gamma\downarrow 0$, the next token will be the token that corresponds to the highest logit value; this is called greedy decoding. We note that the greedy decoding is deterministic, i.e., the token is sampled from a degenerate distribution. These two families of decoding methods can be unified as Algorithm 1, where LLMs autoregressively generate tokens until the $\mathrm{EOS}$ token.
@@
Speculative Decoding As shown in Algorithm 1, the autoregressive decoding feature requires multiple forward inferences of LLMs $P$ sequentially. To reduce the number of forward inferences, leviathan2023fast; chen2023accelerating proposed the vanilla speculative decoding algorithm, which implements a draft model $Q$ to generate draft tokens and let the target model $P$ verify them in parallel. For completeness, we present and describe the vanilla speculative decoding algorithm in Appendix C.1. This vanilla speculative decoding is then extended by some existing works, e.g., miao2023SpecInferAG; cai2024medusa organises the draft tokens as a tree, which improves the number of accepted tokens. The speculative decoding algorithm contains many hyperparameters, e.g., the draft model $Q$, and the tree structure in miao2023SpecInferAG; cai2024medusa. Most existing works keep these hyperparameters fixed for all the tasks. Some other works optimise the draft model in an online or offline manner (liu2023OnlineSD) and the size of the tree (chen2024SequoiaSR), which are designed for specific considerations. In contrast, our work aims to derive a unified online hyperparameter selection algorithm that can be applied for any type of hyperparameters.
Multi-Armed Bandits The Multi-Armed Bandit (MAB) is a fundamental online decision-making problem (see Algorithm 7 for its dynamics). In its classical stochastic form, an agent chooses from $K$ arms, each of which delivers a reward sampled i.i.d. from an unknown but fixed distribution when pulled (lattimore2020bandit). The goal is to select arms over $T$ rounds to maximise cumulative rewards. Two primary classes of algorithms—UCB-type (auer2002finite) and sampling-based methods (russo2017ATO)—have been developed and proven optimal in this setting. In the adversarial formulation, there are no assumptions on the reward distributions; rewards can evolve arbitrarily over time and may be correlated across arms (auer2002nonstochastic). Several algorithms, such as EXP3 and EXP4 (auer2002nonstochastic), are known to achieve optimal performance under these conditions. In this work, we frame the hyperparameter selection problem as an MAB problem and develop algorithms tailored to both stochastic and adversarial settings.
Notations: Let $N]:=\{1,\cdots,N\}$. For a finite set $\mathcal{X}$, we denote the set of distributions supported on it as $\bm{\Delta}_{\mathcal{X}}=\{P:\mathcal{X}\rightarrow[0,1]\,|\,\sum_{x\in\mathcal{X}}P(x)=1,P(x)\geq 0\text{ for all }x\in\mathcal{X}\}$. The space of all finite-length sequences whose components belong to $\mathcal{X}$ is denoted as $\mathcal{X}^{*}$, and we use $x_{1:L}\in\mathcal{X}^{L}\subseteq\mathcal{X}^{*}$ to denote a length-$L$ sequence. The Kullback–Leibler ($\mathrm{KL}$) divergence between two distributions $P$ and $Q$ is denoted as $\mathrm{KL}(P,Q)$.
@@ 3 Bandits for Adaptive Speculative Decoding
Algorithm 2 Speculative Decoding Subroutine (SpecDecSub)
@@
To put the problem in a mathematically sound way, we first specify a general speculative decoding subroutine (SpecDecSub) in Algorithm 2. The input of this subroutine is a prompt $\mathrm{pt}\in\mathcal{X}^{*}$, a target model $P$, a specification of hyperparameters $S$, and the maximum speculation length $L$, and the output is the accepted token sequence $x_{1:\tau}\in\mathcal{X}^{*}$. We provide two examples of the hyperparameter sets here. (1) If we adopt the vanilla speculative decoding (Algorithm 6) as Line 1, $S$ can be different draft models $Q:\mathcal{X}^{*}\rightarrow\bm{\Delta}_{\mathcal{X}}$, and $\mathcal{S}$ is the set of all the provided draft models. We would like to choose a draft model according to its training context, e.g. math, creative writing, to decode the current prefix. Then the problem we consider is how to adaptively select a proper draft model for speculative decoding via bandit algorithms. (2) If we adopt Medusa (cai2024medusa) as Line 1, $S$ can be different tree structures, and $\mathcal{S}$ is the set of plausible tree structures. In this problem, we would like to adaptively adjust the speculation tree structure according to the context.
@@
Before going to the algorithm design and the theoretical analysis, we would like to specify some important properties that are shared by any arm selection algorithm $\mathtt{ALG}$ and clarify the intuitions about our theoretical analysis. We denote the stopping time of the canonical decoding (Algorithm 1) and the generated sequence as $\tau_{\mathrm{c}}$ and $\mathrm{pt}_{\tau_{\mathrm{c}}}$, respectively.
@@
$\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})}\overset{d}{=}\mathrm{pt}_{\tau_{\mathrm{c}}},\text{ and }\mathrm{len}(\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})})\overset{d}{=}\mathrm{len}(\mathrm{pt}_{\tau_{\mathrm{c}}}).$ (3)
@@
Interpretation of the Desired Result. Given a prompt $\mathrm{pt}$ and bandit configuration $\nu$, BanditSpec adaptively selects the hyperparameter via $\mathtt{ALG}$ and learns the context. The stopping time regret (1) measures how the stopping time of $\textsc{BanditSpec}(\mathtt{ALG})$ compares to that of the (agnostic) best one $\textsc{BanditSpec}(\mathtt{ALG}_{i^{*}})$. By minimizing $\mathrm{Reg}(\mathtt{ALG},\mathrm{pt},\nu)$, we want to devise an $\mathtt{ALG}$ to (approximately) match the performance of $\mathtt{ALG}_{i^{*}}$. In particular, if $\mathrm{Reg}(\mathtt{ALG},\mathrm{pt},\nu)=o(\mathbb{E}[\mathrm{len}(\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})})])$, it implies that $\textsc{BanditSpec}(\mathtt{ALG})$ requires the same number of speculative decoding rounds as $\textsc{BanditSpec}(\mathtt{ALG}_{i^{*}})$ asymptotically even though the information about $S_{i^{*}}$ is not revealed at the beginning. In other words, $\textsc{BanditSpec}(\mathtt{ALG})$ learns the identity of $S_{i^{*}}$ quickly and the price for this learning process can be amortized over time. Additionally, when $\textsc{BanditSpec}(\mathtt{ALG})$ is deployed over diverse prompt inputs, we expect a significant acceleration of token generation compared to any fixed single speculative decoding method.
@@ 4 Modeling Tokens Stochastically
Figure 2: Illustration of our bandit model for choosing configurations to decode the next token, where UCB and EXP3 refer to UCBSpec and EXP3Spec, respectively.
@@ 4.1 Upper Bounds for the Stochastic Case
Algorithm Design We design a UCB-type arm selection algorithm UCBSpec, as shown in Algorithm 4. To avoid additional terms, we call the aggregated algorithm, $\textsc{BanditSpec}(\textsc{UCBSpec})$, as UCBSpec. The full version of UCBSpec is detailed in Algorithm 8.
This aggregated algorithm, UCBSpec, is adapted from the classical UCB-1 algorithm in auer2002finite. The main differences are the confidence radius design $\mathrm{cr}_{i,t}$ and the stopping rule. We highlight that the form of $\mathrm{cr}_{i,t}$ is designed to fit the weak assumption of the number of accepted tokens. In fact, the proof of the regret of UCB-1 assumes that the values of each arm are generated before the pull of arms (auer2002finite; lattimore2020bandit), which bifurcates from practical LLM inference scenarios. In contrast, we remove this strong restriction. The stopping rule of UCBSpec makes the analysis of our algorithm rather different from that of UCB-1. The stochasticity of the total number of arm pulls requires a novel regret decomposition analysis that is not presented in previous works.
@@ 4.2 Lower Bound for the Stochastic Case
This TGD was considered in the seminal works on speculative decoding (leviathan2023fast; chen2023accelerating).
@@
For the tightness of our algorithm, according to Note 15.3 in lattimore2020bandit, $\mathrm{kl}_{i}=O(\Delta_{i}^{2})$ when $\Delta_{i}$ is small. This indicates the dominating terms in the upper bound in Theorem 4.3 match the lower bound in Theorem 4.4 up to (possibly instance-dependent) constants. Furthermore, because the truncated geometric distribution is closer to a sub-exponential family distribution, especially when $L$ is large, bandit algorithms built upon UCB1 (auer2002finite) are generally loose in some factors. In order to close the gap between the upper and lower bounds completely, KL-UCB (garivier2011KLUCB) can possibly be adapted to this problem out of theoretical interest. However, on the practical side, KL-UCB demands solving an optimization problem at each round, which can be time-consuming during implementations. Thus, it does not perfectly align with our ultimate goal of LLM inference acceleration.